Q=x1^2+3x2^2+3x3^2-2x2x3
1
Expert's answer
2021-09-21T00:27:36-0400

"Q=(x_1)^2+3(x_2)^2+3(x_3)^2-2(x_2)(x_3)\\
Q=(x_1)^2+2(x_2)^2+2(x_3)^2+(x_2)^2+(x_3)^2-2(x_2)(x_3)\\
Q=(x_1)^2+2(x_2)^2+2(x_3)^2+(x_2-x_3)^2\\
Q=

\begin{bmatrix}
 x_1 \\
 x_2 \\
x_3 
\end{bmatrix}
\begin{bmatrix}
 x_1 & 2x_2 & 2x_3\\
 
\end{bmatrix} +\begin{bmatrix}
 x_2-x_3 
\end{bmatrix}\begin{bmatrix}
 x_2-x_3 
\end{bmatrix}"